Fermi National Laboratory
 

Accelerator Update for July 4, 2007 - July 6, 2007

 

Wednesday July 4

The day shift began with Operations monitoring the Tevatron (TeV) store 5533 (101.39E30) and the Recycle stash (95.5E10), with the Antiproton Source (Pbar) stacking (16.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson MTest area taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 9:53 AM, the MI beam aborted and the MI beam permit went down when both abort kickers tripped off on a “prefire/fail to fire” indication (I:KPS4A & B).

At 3:04 PM, the current on a Meson power supply (F:M01D) drifted low, causing MTest not to receive beam. Operators adjusted the supply’s current.

Operations reported at 4:06 PM that the Recycler had stashed antiprotons three times during the day shift.

At 4:32 PM, two MI kickers (I:KPS4A & B) tripped off. They reset.

Operations disabled the MTest beam at 6 PM.

At 9:56 PM, the MI Power Supply interlock loop tripped off on a ground fault. Eventually, everything was reset. Beam resumed at 10:36 PM.

Operations terminated store 5533 at 11 PM.

Operations prepared D0 for an access at 11:16 PM.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Thursday July 5

The midnight shift began with D0 in access, with Operations monitoring the stash (299.7E10), with Pbar stacking (12.9E10), and with NuMI and MiniBooNE taking beam.

D0 came out of access at 1:45 AM.

At 1:46 AM, KPS4A & B tripped off. This was followed by a vacuum pump trip (IP7144) and a vacuum burst. Beam returned at 1:58 AM.

Operations put the TeV through a dry squeeze at 2:07 AM.

Operations put the TeV into shot setup at 2:31 AM.

At 2:56 AM, the Recycler lost about 6 mA of the stash due to a damper problem. Also, the emittances got large. Operators contacted experts. This put shot setup on hold.

At 4:19 AM, Operations halted the TeV shot setup and put the TeV into a dry squeeze.

Operations prepared CDF for an access at 5:36 AM to recover their silicon detector.

At 5:56 AM, MTest requested beam. Operations tried to send them beam, but found that M01D was off. MTest began receiving beam at 6:35 AM.

At 6:10 AM, the Recycler suffered from an emittance blowup.

At 6:15 AM, KPS4A & B tripped off for the first of two times in 75 minutes. They were reset.

At 8:48 AM, a NuMI power supply (HV101) tripped off on a ground fault indication. A power supply expert investigated, but the alarm had cleared by the time he got there. NuMI resumed taking beam at 9:21 AM.

At 9:26 AM, a Linac Klystron RF station (KRF1) tripped off. It reset.

The Neutron Therapy Facility reported at 10:56 AM that they had treated the last patient for the day.

TeV experts began a two-hour study at 10:04 AM.

At 12:46 PM, KPS4A & B tripped off. They were reset.

Operations prepared CDF and D0 for an access at 1:12 PM.

At 1:39 PM, a Linac expert reported that LRF3 was sparking and it required immediate attention. Operations halted the Linac beam for this work.

At 2:39 PM, the Meson M01 interlocks dropped. Operations contacted the interlock technicians. They investigated and repaired some loose wires. MTest was down for a little less than two hours.

A Linac expert reported that they were through with their work on LRF3. Linac beam resumed at 2:44 PM.

A MI expert began a proton coalescing study at MI-60 at 3:06 PM.

CDF came out of access at 3:46 PM.

A Switchyard (SY) expert reported at 3:49 PM that MTest would not take beam on Friday for a scheduled maintenance day.

Operations reported at 4:02 PM that the Recycler had stashed once during the day shift.

D0 came out of access at 4:54 PM.

Operations put the TeV through a wet squeeze at 5:02 PM.

At 5:18 PM, the MI abort kickers again failed. Operators noted the status and then reset them.

TeV experts began a study period at 5:30 PM.

Operations disabled the MTest experiment at 6:08 PM.

At 6:54 PM, the TeV suffered a quench at sectors DB and C4 during their studies. A TeV expert investigated and placed the blame on a power supply (T:VA18). Operations changed out a control card.

At 7:06 PM, KPS4A & B tripped off for the first of three times in about 40 minutes. It was reset each time.

At 9:46 PM, Operations reported that the H- Source’s output was dropping. An expert was called.

At 10:04 PM, the NuMI HV101 power supply tripped off on another ground fault. An expert was called in. Operations halted MI beam for this investigation.

Operations put the TeV into shot setup at 11:08 PM.

The power supply expert changed out the NuMI HV101 ground fault circuitry. NuMI began receiving beam at 11:09 PM.

Operations reported at 11:56 PM that the Recycler had stashed once during the evening shift.

Friday July 6

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (201.8E30), with Pbar stacking (111.5E10), and with NuMI and MiniBooNE taking beam.

Operations established store 5540 at 12:50 AM with an initial luminosity of 191.78E30.

At 12:57 AM, the Linac low energy buncher tripped off for the first of two times in about four hours. It reset.

A Linac expert reported at 2:29 AM that he had tuned up the H- Source as best he could, but at a level far down from what it was running at on Thursday.

Operations made the Meson M02 – 5 enclosures ready for access at 5:42 AM.

Operations reported at 7:57 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow a MI slipstacking study. Next week Wednesday at 6 AM, 7/11/07, CDF needs to make a two-shift access to make necessary repairs.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/06/2007    email Fermilab